# TL-Training: A Task-Feature-Based Framework for Training Large Language Models in Tool Use

### v1

@@ 6 Further Studies
To further demonstrate the strengths of our approach, we conduct additional analysis focusing on both robustness (Section 6.1) and general performance (Section 6.2).
@@ 6.2 General Performance
The strong performance of LLMs is largely attributed to their extensive world knowledge and generalizability acquired during pre-training (Ye et al., 2023). However, fine-tuning on domain-specific tasks can sometimes compromise this generalizability (Yang et al., 2024b; Ghosh et al., 2024). To assess whether TL-CodeLLaMA-2’s general-purpose capabilities are affected by its exclusive training on tool-use data, we evaluate its performance on three general test sets: MMLU (knowledge) (Hendrycks et al., 2021), GSM8K (math) (Cobbe et al., 2021), and HumanEval (code) (Chen et al., 2021), comparing it to CodeLLaMA-2-7B. As shown in Figure 5, despite being fine-tuned solely for the tool-use task, TL-CodeLLaMA-2 retains its original knowledge and task performance and even shows slight improvements in math and coding abilities. This may be because our method requires only a small amount of training data, resulting in minimal changes to the model’s original parameters, thus preserving its knowledge base (Ren et al., 2024; Wang et al., 2024; Ye et al., 2024d). Furthermore, the enhancement in tool-use ability appears to improve the model’s reasoning capacity, contributing to better performance in math and coding tasks. These findings further underscore the broad applicability of our approach.
7 Related Works
Training LLMs for Tool Use
LLMs capable of utilizing external tools significantly enhance their ability to interact with dynamic environments and address diverse user needs (Qin et al., 2023). However, the diversity and complexity of real-world tools present significant challenges in training such models. Existing methods, such as SFT, rely on the generation of extensive datasets of tool interactions (Song et al., 2023; Tang et al., 2023), enabling models to learn tool functionalities, invoke appropriate tools, and process feedback effectively. While effective, these methods are resource-intensive due to the large-scale data construction involved. To overcome these challenges, some studies have proposed encoding tool names as special tokens directly integrated into model training, embedding tool-specific knowledge into the model’s intrinsic capabilities (Hao et al., 2023). This approach has shown promise for existing tools but remains limited in its ability to adapt to newly introduced tools. Building on these findings, our work introduces a novel training paradigm for tool-use LLMs, addressing both efficiency and adaptability. By leveraging a compact dataset of 1,217 data points and incorporating three task-specific components, our approach achieves state-of-the-art performance while significantly reducing data requirements.
Evaluating LLMs in Tool Use
Evaluating LLMs’ tool-use capabilities is essential for understanding their effectiveness in diverse scenarios. A common evaluation method involves comparing predicted outputs with standard answers from a single turn of tool usage (Chen et al., 2024). However, in multi-turn interactions, the variability in invocation processes complicates the definition of a single standard path. To address this, evaluations increasingly consider multiple dimensions of tool-use processes and outcomes (Ye et al., 2024a). Beyond tool-use performance, researchers have also investigated robustness and safety in practical scenarios. For example, robustness benchmarks like RoTBench (Ye et al., 2024c) and safety-focused studies such as ToolSword (Ye et al., 2024b) provide insights into how LLMs manage edge cases and avoid harmful outputs. In this paper, we evaluate LLMs across single-turn and multi-turn tool use to provide a more comprehensive assessment. Additionally, we analyze robustness to further demonstrate the superiority of our approach under diverse conditions.
@@ Appendix A Theorems and Proofs
Appendix B Prompt Template
We use the official prompt of each LLM for tool use, with corresponding examples provided from Table 6 to Table 13.
@@ Appendix B Prompt Template
You are AutoGPT, you can use many tools(functions) to do the following task. First I will give you the task description, and your task start. At each step, you need to give your thought to analyze the status now and what to do next, with a function call to actually excute your step. Your output should follow this format: Thought: Action Action Input: After the call, you will get the call result, and you are now in a new state. Then you will analyze your status now, then decide what to do next… After many (Thought-call) pairs, you finally perform the task, then you can give your finial answer. Remember: 1.the state change is irreversible, you can’t go back to one of the former state, if you want to restart the task, say Ï give up and restart.̈ 2.All the thought is short, at most in 5 sentence. 3.You can do more then one trys, so if your plan is to continusly try some conditions, you can do one of the conditions per try. Let’s Begin! Task description: You should use functions to help handle the real time user querys. Remember: 1.ALWAYS call F̈inishf̈unction at the end of the task. And the final answer should contain enough information to show to the user,If you can’t handle the task, or you find that function calls always fail(the function is not valid now), use function Finish-¿give_up_and_restart. 2.Do not use origin tool names, use only subfunctions’ names. Specifically, you have access to the following APIs: [{“type”: “function”, “function”: {“name”: “random_advice”, “description”: “Returns a random advice slip as a slip object.”, “parameters”: {“type”: “object”, “properties”: {}, “required”: []}}}] User: Can you fetch some random advice for me? Assistant: |
@@
你是一个名为 ChatGLM 的人工智能助手。你是基于智谱AI训练的语言模型 GLM-4 模型开发的,你的任务是针对用户的问题和要求提供适当的答复和支持。 # 可用工具 ## random_advice {“name”: “random_advice”, “description”: “Returns a random advice slip as a slip object.”, “parameters”: {“type”: “object”, “properties”: {}, “required”: []}} 在调用上述函数时,请使用 Json 格式表示调用的参数。user Can you fetch some random advice for me?assistant |
@@
Appendix C Examples of Tool Call and Feedback
In Table 14, we present examples of potential scenarios that may arise during tool calls to provide readers with a clearer and more visual understanding.
@@ Appendix C Examples of Tool Call and Feedback
| Tool Instability | GPT-4-turbo | {“arguments”: “{“apikey”: “your_api_key_here”}”, “name”: “get_cryptocurrency_categories”} | HTTPSConnectionPool(host=…, port=443): Max retries exceeded with url: … |
@@
The function cannot be executed successfully with the given parameters for the reason: 400 Client Error: Bad Request for url: … |
| Tool Instability | Qwen-2-Instruct |
Thought: …
Action: search_word Action Input: {“api_key”: “”, “letterPattern”: “a*b*c”} |
{“error”: “connection error.”} |
@@
| Tool Call Failure | GPT-3.5-turbo | {“arguments”: “{“query”: “Bitcoin”, “data_type”: “RELATED_QUERIES”, “limit”: “5”}”, “name”: “google_trends_search”} | ‘¡=’ not supported between instances of ‘str’ and ‘int’ |
@@
The function cannot be executed successfully with the given parameters for the reason: Object of type bytes is not JSON serializable |
| Tool Call Failure | GPT-3.5-turbo | {“arguments”: “{“destination”: “Seattle”, “exhibit_maxnum”: “3”}”, “name”: “lodgingProducts”} | slice indices must be integers or None or have an __index__ method |
@@
Action: get_news_headlines Action Input: { “api_key”: “your_api_key”, “q”: “technology”, “sortBy”: “popularity”, “pageSize”: “5” } |
The function cannot be executed successfully with the given parameters for the reason: get_news_headlines() got an unexpected keyword argument ‘sortBy’ |
| Parameter Missing | Qwen-2.5-Instruct | {“arguments”: “{“q”: “London”, “days”: “7”, “api_key”: “your_api_key”}”, “name”: “forecast”} | forecast missing 1 required positional argument: aqi |
| Others | GPT-4o | {“arguments”: “{“website”: “www.mywebsite.com”}”, “name”: “analyze_scan”} | {“error”: “recent-scan-not-found”, “text”: “Recently completed scan for www.mywebsite.com not found”} |

### v2

@@ 7.2 General Performance
The strong performance of LLMs is largely attributed to their extensive world knowledge and generalizability acquired during pre-training Ye et al. (2023). However, fine-tuning on domain-specific tasks can sometimes compromise this generalizability Yang et al. (2024c); Ghosh et al. (2024). To assess whether TL-CodeLLaMA-2’s general-purpose capabilities are affected by its exclusive training on tool-use data, we evaluate its performance on three general test sets: MMLU (knowledge) Hendrycks et al. (2021), GSM8K (math) Cobbe et al. (2021), and HumanEval (code) Chen et al. (2021), comparing it to CodeLLaMA-2-7B. As shown in Figure 5, despite being fine-tuned solely for the tool-use task, TL-CodeLLaMA-2 retains its original task performance and even shows slight improvements in math and coding abilities. This is because our method requires only a small amount of training data, resulting in minimal changes to the model’s original parameters, thus preserving its knowledge base Ren et al. (2024); Wang et al. (2024); Ye et al. (2024c). Furthermore, the enhancement in tool-use ability appears to improve the model’s reasoning capacity, contributing to better performance in math and coding tasks. These findings further underscore the broad applicability of our approach.
@@ 8 Conclusion
Acknowledgments
The authors wish to thank the anonymous reviewers for their helpful comments. This work was partially funded by National Natural Science Foundation of China (No.62476061,62206057), Shanghai Rising-Star Program (23QA1400200), Natural Science Foundation of Shanghai (23ZR1403500).
@@ Appendix A Theorems and Proofs
Appendix B Details of Datasets
As shown in Table 2, we construct a custom training set focused on multi-turn tool use and evaluate it using four publicly available test sets.
Training Data
While previous studies often construct over 100,000 data points for training Qin et al. (2024), we deliberately limit our dataset size. Our main goal is to validate the effectiveness of our approach rather than to scale data volume. Surprisingly, the experimental results in Section 6 show that training on just 1,217 data points using our method matches or even exceeds the performance of leading LLMs.
Test Sets
To comprehensively evaluate LLM tool-use performance, we use four open-source tool usage test sets. ToolAlpaca (eval_real), RoTBench (Clean), and BFCL-v3 (executable) are selected for single-turn tool use evaluations, while ToolEyes is used for multi-turn tool use assessment.
Appendix C Details of Baselines
Tool-Use LLMs
ToolLLaMA-2-7B and NexusRaven-2-13B are prominent tool-use LLMs, built on LLaMA-2-7B and CodeLLaMA-2-13B, respectively. These models are trained on a large volume of tool-use data, enhancing their ability to interact with tools. For example, ToolLLaMA-2-7B is trained on over 120,000 data points covering more than 16,000 tools using standard SFT, significantly boosting its tool-use capabilities.
Open-Source LLMs
Among the existing open-source, general-purpose LLMs, ChatGLM-4-chat-9B, Qwen-2-Instruct-7B, LLaMA-3.1-Instruct-8B and Qwen-2.5-Instruct-7B have been specifically optimized for tool use, enabling them to interact with various tools to fulfill user needs.
Closed-Source LLMs
Our Model
Appendix D Prompt Template
@@
You are AutoGPT, you can use many tools(functions) to do the following task. First I will give you the task description, and your task start. At each step, you need to give your thought to analyze the status now and what to do next, with a function call to actually excute your step. Your output should follow this format: Thought: Action Action Input: After the call, you will get the call result, and you are now in a new state. Then you will analyze your status now, then decide what to do next… After many (Thought-call) pairs, you finally perform the task, then you can give your finial answer. Remember: 1.the state change is irreversible, you can’t go back to one of the former state, if you want to restart the task, say Ï give up and restart.̈ 2.All the thought is short, at most in 5 sentence. 3.You can do more then one trys, so if your plan is to continusly try some conditions, you can do one of the conditions per try. Let’s Begin! Task description: You should use functions to help handle the real time user querys. Remember: 1.ALWAYS call F̈inishf̈unction at the end of the task. And the final answer should contain enough information to show to the user,If you can’t handle the task, or you find that function calls always fail(the function is not valid now), use function Finish->give_up_and_restart. 2.Do not use origin tool names, use only subfunctions’ names. Specifically, you have access to the following APIs: [{“type": “function", “function": {“name": “random_advice", “description": “Returns a random advice slip as a slip object.", “parameters": {“type": “object", “properties": {}, “required": []}}}] User: Can you fetch some random advice for me? Assistant: |
@@
你是一个名为 ChatGLM 的人工智能助手。你是基于智谱AI训练的语言模型 GLM-4 模型开发的,你的任务是针对用户的问题和要求提供适当的答复和支持。 # 可用工具 ## random_advice {“name": “random_advice", “description": “Returns a random advice slip as a slip object.", “parameters": {“type": “object", “properties": {}, “required": []}} 在调用上述函数时,请使用 Json 格式表示调用的参数。user Can you fetch some random advice for me?assistant |
@@
Appendix E Examples of Tool Call and Feedback
In Table 14, we present some examples of potential scenarios that may arise during tool calls.
@@ Appendix E Examples of Tool Call and Feedback
| Tool Instability | GPT-4-turbo | {“arguments”: “{}”, “name”: “get_currency”} | HTTPSConnectionPool(host=…, port=443): Max retries exceeded with url: … |
@@
400 Client Error: Bad Request for url: … |
@@
| Tool Call Failure | GPT-3.5-turbo | {“arguments”: “{“query”: “Bitcoin”, “data_type”: “RELATED_QUERIES”, “limit”: “5”}”, “name”: “google_trends_search”} | ‘’ not supported between instances of ‘str’ and ‘int’ |
@@
Object of type bytes is not JSON serializable |
@@
Action: get_news_headlines Action Input: { “api_key”: “your_api_key”, “q”: “technology”, “sortBy”: “popularity”} |
get_news_headlines() got an unexpected keyword argument ‘sortBy’ |
| Parameter Missing | Qwen-2.5-Instruct | {“arguments”: “{“q”: “London”, “days”: “7”, “api_key”: “your_api_key”}”, “name”: “forecast”} | forecast() missing 1 required positional argument: aqi |
| Others | GPT-4o | {“arguments”: “{“website”: “www.mywebsite.com”}”, “name”: “analyze_scan”} | {“error”: “Recently completed scan for www.mywebsite.com not found”} |